OpenStudy (anonymous):

Is this correct? Find the product. -3z(z^3-3z^2+6) =-3z(z^3)+(-3z)(-3z^2)-3z(6) =-3z^4+9z^3-18z =6z-18z =-12z

jimthompson5910 (jim_thompson5910):

-3z(z^3-3z^2+6) =-3z(z^3)+(-3z)(-3z^2)-3z(6) ... Correct =-3z^4+9z^3-18z .... Correct =6z-18z ... Incorrect. There are no like terms to combine. =-12z ... Again there are no like terms to combine.

OpenStudy (anonymous):

so then the answer would be 6z-18z ?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

-3z^4+9z^3-18z is the correct answer

jimthompson5910 (jim_thompson5910):

everything after that is incorrect

OpenStudy (anonymous):

okay, thank you!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (unklerhaukus):

no, mellaflamme86 like jim_thompson5910 says \[-3z(z^3-3z^2+6)=...=-3z^4+9z^3-18z \] and you cannot simplify further